0070171391:

Wyatt's dog is tethered to a post in his yard. If the rope tied to the dog is 12 feet long, how many square yards does the dog have access to? (Use 3.14 for π.)

theycraveekiki13:

CAN YOU HELP ME WITH MY MATH PROBLEM

0070171391:

probably

theycraveekiki13:

OK THANK YOU'

Shadow:

 Solve for the area of a circle \[A = \pi r^2\]

Shadow:

\[A = \pi (12)^2 \] \[A = (3.14)(12)^2 \]

0070171391:

452.16 yd 2 50.24 yd 2 75.36 yd 2 25.12 yd 2

Shadow:

Look at the steps I have shown you above. Can you figure the answer out for yourself?

0070171391:

i will try

Shadow:

What did you get?

0070171391:

452.16 yd 2

Shadow:

That's what I got as well

0070171391:

thank you